OpenStudy (anonymous):

Help Pleaseee!!!! Solve 3^x – 4 = 7x + 9

OpenStudy (anonymous):

\[3^{x-4}=7^{x+9}\]

OpenStudy (anonymous):

3^x-7x=13 or 3^x=7x+13\[3^x-7x=13 OR 3^x=7x+13 \]

OpenStudy (anonymous):

hope you like my answer

OpenStudy (anonymous):

@bahrom7893 Bah explain Lesha how to solve this

OpenStudy (anonymous):

i dont understand it :(

OpenStudy (anonymous):

i got the answer you do explanation

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

one second Lesha bahrom7893 will explain he is pro on them

OpenStudy (bahrom7893):

lol what are we doing again? I just got back from work.

OpenStudy (bahrom7893):

\[3^{x – 4} = 7^{x + 9}\]\[3^x/3^4 = 7^x * 7^9\]

OpenStudy (anonymous):

Solving Exponential Equations with Unequal Bases

OpenStudy (bahrom7893):

\[3^x/7^x=7^9*3^4\]

OpenStudy (anonymous):

We have : \[3^{x-4}=7^{x+9}\] By taking the natural logarithm of the two sides : \[\ln 3^{x-4}=\ln 7^{x+9}\] So : \[(x-4)\ln 3=(x+9)\ln 7\] So : \[x\ln 3-4\ln 3=x\ln7+9\ln7\] So : \[x\ln3-x\ln7=9\ln7+4\ln3\] so : \[(\ln 3-\ln7)x=9\ln7+4\ln3\] So : \[x=\frac{9\ln7+4\ln3}{\ln3-\ln7}\]

OpenStudy (bahrom7893):

oh man you could use lns... nevermind me, the heat wave is having adverse effects on me.

OpenStudy (anonymous):

sooo how would i finish it to find the answer ?

OpenStudy (bahrom7893):

he gave you the answer.

OpenStudy (anonymous):

x ≈ –25.86 x ≈ –2.59 x ≈ –0.39 x ≈ –0.04 are the options ..

OpenStudy (anonymous):

@bahrom7893 take look their is other opetions

OpenStudy (anonymous):

By using the calculator you can get the answer : \[\approx -25.86\]

OpenStudy (anonymous):

thnkyou (:. can you help with one more ?

OpenStudy (anonymous):

Solve 5^x + 5 = 9^x

OpenStudy (anonymous):

Are you sure that the equation is like this : \[5^x+5=9^x\]??

OpenStudy (anonymous):

\[5^{x+5}=9^{x}\]

OpenStudy (anonymous):

yes thats how it looks..^^

OpenStudy (anonymous):

By the same method that I solved the previous exercise ! By taking the natural logarithm of the two sides : \[\ln5^{x+5}=\ln9^x\] So : \[(x+5)\ln5=x\ln9\] so : \[x\ln5+5\ln5=x\ln9\] so : \[x\ln5-x\ln9=-5\ln5\] So : \[(\ln5-\ln9)x=-5\ln5\] So : \[x=\frac{5\ln5}{\ln9-\ln5}\approx 13.69\]

OpenStudy (anonymous):

Omgg ii Got A Hundred ! Thankyouu Sooo Muchhhh (;: !

OpenStudy (anonymous):

@Lesha You are welcome